# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v14

@@ 2.3 Statistical Taylor Expansion
$\zeta(n)\equiv$ $\int_{\mu-\varrho\sigma}^{\mu+\kappa\sigma}\tilde{x}^{n}\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
@@ 2.4 Bounding Symmetry
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion. Both Gaussian and Uniform distributions are symmetric.
2.5 Bounding Asymptote
Empirically, Formula (2.18), (2.22), and (2.25) demonstrate that as $2n\rightarrow+\infty$, $\zeta(2n)\rightarrow\kappa^{2n}/(2n)$.
2.5.1 Uniform Input Uncertainty
For uniform distribution, the bounding range is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.18) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
$\zeta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.18)
2.5.2 Gaussian Input Uncertainty
Formula (2.7) reduces to Formula (2.19) and (2.20):
$\zeta(2n,\kappa)=$ $\;(2n-1)!!\left(\xi(\frac{\kappa}{\sqrt{2}})-2N(\kappa)\sum_{j=0}^{n-1}\frac{\kappa^{2j+1}}{(2j+1)!!}\right);$ (2.19)
$=$  $\;2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.20)
$=$ $\;(2n-1)\zeta(2n-2,\kappa)-2N(\kappa)\kappa^{2n-1};$ (2.21)
$\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)\simeq 2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.22)
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.21). When $\kappa=5$, and $n<5$, the relative error $\left|\zeta(2n)/(2n-1)!!-1\right|$ is less than $10^{-3}$.
  • For large $2n$, Formula (2.20) reduces to Formula (2.22), showing that $\zeta(2n)$ increases more slowly than $\kappa^{2n}$ as $2n$ grows.
2.5.3 An Input Uncertainty with Limited Range
$\rho(\tilde{x},\mu,\sigma)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;\mu=2\lambda;\;\;\;\sigma=\sqrt{2}\lambda;$ (2.23)
$e^{-\varrho}\varrho^{2}$ $=e^{-\kappa}\kappa^{2};$ (2.24)
$\lim_{n\rightarrow+\infty}\zeta(n,\kappa)$ $=\frac{\frac{\kappa}{\sqrt{2}}^{n}}{n+2}e^{-\kappa};$ (2.25)
Formula (2.23) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<2<\kappa$. Formula (2.24) gives its mean preserving equation. And Formula (2.25) describes its asymptotic behavior.
@@ 2.7 Low-Order Approximation
2.8 Convergence
@@ 2.10 Ideal Statistics
In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. Define the ideal ratio to be zero when the sample count $N$ is $1$. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
@@ 3.3 Monotonic
When the input uncertainty is Uniformly distributed, Figure 10 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.46). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 9 and 10 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 9 are $3$ order-of-magnitude smaller than those in Figure 10 when the exponent $c$ is less than 0. Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
@@ 9.9 Using Library Sine for Sin/Cos Signals
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 40 and 46.
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 50. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 45. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 49. Also, it is completely absent when using the Quart sine function, as seen in Figure 38 and 39. Figure 50 and Figure 51 demonstrate that the error deviations increase with sine or cosine frequency. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 44, 45, and 48. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.12 Using Library Sine for Linear Signal
As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional unspecified errors.
@@ 9.13 Ideal Coverage
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 58 and 59 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 60 and 61 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region. Figure 61 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 37, so that the error deviations within the proper coverage region differ from $1$ only marginally.
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 60 and 61 are almost identical using either C++, or Java, or Python.
@@ 9.14 Input Uncertainty Recovery
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 40, 42, 46, and 52, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 54. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized. Perhaps the asymptotic behavior implies a generic new requirements on the probability distributions.

### v15

@@ 2.3 Statistical Taylor Expansion
$\zeta(n)\equiv$ $\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
@@ 2.4 Bound Moment Symmetry
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
@@ 2.6 Low-Order Approximation
2.7 Asymptotic Behavior for Bound Moment
Formula (2.38) is a special case of Formula (2.37) for symmetric probability distribution. Empirically, Formula (2.38) is confirmed by Formula (2.40), (2.42), and (2.43), and Formula (2.37) is confirmed by Formula (2.45) and (2.51).
$\kappa\ll n$ $:\;\;\;\zeta(n,\kappa)\simeq\rho(\kappa)\frac{\kappa^{n}}{n};$ (2.37)
[…]
2.7.1 Gaussian Input Uncertainty
Figure 5: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus Formula (2.41) (left y-axis), and their ratio (right y-axis).
With Formula (2.39) [Double_Factorial], for Gaussian distribution, Formula (2.40) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, while Formula (2.41) demonstrates its asymptotic behavior. Figure 5 shows the transition to Formula (2.41) with increasing order $2n$, in which the ratio is the ratio of Formula (2.41) when $\kappa=5$ to $\zeta(2n,5)$.
$\sum_{j=0}^{+\infty}\frac{\kappa^{2j+1}}{(2j+1)!!}=\sqrt{\frac{\pi}{2}}\xi(\frac{\kappa}{\sqrt{2}})e^{\frac{1}{2}\kappa^{2}};$ (2.39)
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}N(\tilde{x})d\tilde{x}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.40)
$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
2.7.2 Uniform Input Uncertainty
As shown in Figure 38, rounding errors can be uniformly distributed [29].
For uniform distribution, the bounding range $\kappa$ is less than the maximal range $\sqrt{3}$. Formula (2.42) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2\sqrt{3}}d\tilde{x}=2\frac{1}{2\sqrt{3}}\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.42)
2.7.3 Laplace Input Uncertainty
For Laplace distribution [4], Formula (2.43) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, and Formula (2.44) demonstrates its asymptotic behavior.
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2}e^{-|\tilde{x}|}d\tilde{x}=e^{-\kappa}\kappa^{2n}\sum_{\kappa=1}^{\infty}\kappa^{j}\frac{(2n)!}{(2n+j)!};$ (2.43)
$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2\frac{1}{2}e^{-\kappa}\frac{\kappa^{2n+1}}{2n+1};$ (2.44)
2.7.4 Exponential Input Uncertainty
An exponential distribution [4] is no longer symmetric. Formula (2.45) provides $\zeta(2n,\kappa)$, Formula (2.46) shows the equation for the mean preserving condition, and Formula (2.47) demonstrates its asymptotic behavior.
$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}(\tilde{x}-1)^{n}e^{-\tilde{x}}d\tilde{x}$ (2.45)
$=e^{-\kappa}(\kappa-1)^{n}\sum_{j=1}^{\infty}(\kappa-1)^{j}\frac{n!}{(n+j)!}-e^{-\varrho}(\varrho-1)^{n}\sum_{j=1}^{\infty}(\varrho-1)^{j}\frac{n!}{(n+j)!};$
$\varrho e^{-\varrho}=\kappa e^{-\kappa}\Longrightarrow 0<\varrho<1<\kappa;$ (2.46)
$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=e^{-\kappa}\frac{\kappa^{n+1}}{n+1};$ (2.47)
2.7.5 Simplified Gamma Input Uncertainty
The gamma distribution [4] $\rho(\tilde{x})=x^{\alpha-1}e^{-x/\lambda}\lambda^{\alpha}/\Gamma(\alpha),\tilde{x},\alpha,\lambda>0$ is simplified as Formula (2.48), whose mean and deviation are $m+1$ and $\sqrt{m+1}$, respectively. Formula (2.49) provides $\zeta(2n,\kappa)$, Formula (2.50) shows the equation for the mean preserving condition, and Formula (2.51) demonstrates its asymptotic behavior.
[…]
$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}\left(\frac{\tilde{x}-(m+1)}{\sqrt{m+1}}\right)^{n}\rho(\tilde{x})d\tilde{x}$ (2.49)
$e^{-\varrho}\varrho^{m+1}=e^{-\kappa}\kappa^{m+1};\Longrightarrow 0<\varrho<m+1<\kappa;$ (2.50)
$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=\frac{1}{m!}e^{-\kappa}\frac{\kappa^{m+n+1}}{m+n+1};$ (2.51)
2.8 Convergence of Variance
When $n\rightarrow\infty$, $\zeta(j,\kappa)\zeta(n-j,\kappa)$ will be ignored in Formula (2.3) because $\zeta(j,\kappa)\zeta(n-j,\kappa)\ll\zeta(n,\kappa)$. The asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
@@ 2.10 Ideal Statistics
In both Figure 8 and 9, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. Define the ideal ratio to be zero when the sample count $N$ is $1$. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.57) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 8 and 9, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
@@ 3.3 Monotonic
When the input uncertainty is Uniformly distributed, Figure 11 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.53). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 10 and 11 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 10 are $3$ order-of-magnitude smaller than those in Figure 11 when the exponent $c$ is less than 0. Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
@@ 9.9 Using Library Sine for Sin/Cos Signals
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 41 and 47.
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 51. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50. Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40. Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency. Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.12 Using Library Sine for Linear Signal
As shown by the difference between Figure 35 and 36, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional unspecified errors.
@@ 9.13 Ideal Coverage
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 59 and 60 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 61 and 62 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region. Figure 62 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 38, so that the error deviations within the proper coverage region differ from $1$ only marginally.
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 61 and 62 are almost identical using either C++, or Java, or Python.
@@ 9.14 Input Uncertainty Recovery
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 41, 43, 47, and 53, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 55. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The necessary and sufficient conditions for Formula (2.37) need to be found, to extend the applicable types of distributions for the statistical Taylor expansion from the most common ones as enumerated in Section 2. The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.